OpenStudy (anonymous):

What is the line of symmetry for the parabola whose equation is y = x^2 - 12x + 7? 1.) x = -6 2.) x = 6 3.) x = 12

OpenStudy (anonymous):

axis will be at \(x=-\frac{b}{2a}\) which in your case is ...

OpenStudy (primeralph):

x = -b/2a. --12/(2*1)

OpenStudy (anonymous):

Is that -12/(2*1) ?

OpenStudy (primeralph):

It's at -b/2a, so the -12 gets another - sign.

OpenStudy (anonymous):

So would it be -12 or 12?

OpenStudy (primeralph):

What is --12?

OpenStudy (anonymous):

Well, that's what I'm asking you. Is it a negative or a positive 12? I'm unsure, considering there's two negative signs in front of the 12.

OpenStudy (anonymous):

the minus sign in the formula \(-\frac{b}{2a}\) and the minus sign in \(-12\) are both there compute \[-\frac{-12}{2}\]

OpenStudy (anonymous):

I don't know what that means.

OpenStudy (anonymous):

two minus signs make it positive

OpenStudy (anonymous):

THANK YOU! Lol

OpenStudy (anonymous):

just like \(5-(-2)=5+2=7\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

So the answer is B?

OpenStudy (anonymous):

the answer is \(x=6\)